# INR-TH-2025-001

### v2

@@ Abstract
We examine the sterile neutrino dark matter production in the primordial plasma with lepton asymmetry unequally distributed over different neutrino flavors. We argue that with the specific flavor fractions, one can mitigate limits from the Big Bang Nucleosynthesis on the sterile-active neutrino mixing angle and sterile neutrino mass. It happens due to cancellation of the neutrino flavor asymmetries in active neutrino oscillations, which is more efficient in the case of inverse hierarchy of active neutrino masses and does not depend on the value of CP-phase. This finding opens a window of lower sterile-active mixing angles. Likewise, we show that, with lepton asymmetry disappearing from the plasma at certain intermediate stages of the sterile neutrino production, the spectrum of produced neutrinos becomes much colder, which weakens the limits on the model parameter space from observations of cosmic small-scale structures (Ly-$\alpha$ forest, galaxy counts, etc.). This finding reopens the region of lighter sterile neutrinos. The new region may be explored with the next generation of X-ray telescopes searching for the inherent peak signature provided by the dark matter sterile neutrino radiative decays in the Galaxy.
@@
However, $\theta$ may also be limited from below if the only source of sterile neutrino DM production is the sterile-active mixing  [^1] There are many models where other interactions are producing DM sterile neutrinos [11, 12, 13, 14, 15].. It is this case which we study below.
@@
The lepton asymmetry may influence the later cosmological stages, in particular, primordial nucleosynthesis and recombination. Observations of primordial helium abundance favor standard cosmology and severely constrain [16, 17] the asymmetry in electron neutrinos at the moment of neutron freeze-out, which happens at the plasma temperature of about $T_{n}\approx 0.7$ MeV, some time after neutrino decoupling. As we show below, the lower bounds on mixing $\theta$, inferred from this constraint, can be significantly reduced in the case of a specific initial lepton asymmetry pattern, with large individual flavor asymmetries but a total lepton asymmetry close to zero. Then active neutrino oscillations, which start at temperature $T\lesssim 20$ MeV, very effectively cancel all flavor asymmetries in accordance with calculations of Ref. [18]. We find that the cancellation is more efficient the in case of inverse ordering of neutrino masses and CP-phase does not play a noticeable role. This allows the larger flavor lepton asymmetries at sterile neutrino production, hence lower mixing angles, consistent with observed helium abundance and anisotropy of the cosmic microwave background.
@@
The sterile neutrino mass $m_{\nu_{s}}$ is also bounded from below by observations of the cosmic structure formation. Being produced in oscillations of the active neutrino, the sterile neutrino inherits its momentum, which is generically of order of the plasma temperature. Hence, due to redshifting, a keV-scale sterile neutrinos exhibit velocities of order $10^{-3}$ at the onset of cosmic structure formation, at the epoch of equality between radiation and matter, when the plasma temperature is $T\approx 0.7$ eV. The dark matter candidates with this feature are called warm dark matter. The matter inhomogeneities at short distances are smoothed off by the free streaming of such dark matter particles. Observations of dwarf galaxies constrain the warm dark matter models with too high velocities, in particular, exclude too light sterile neutrinos, for recent studies see e.g. [19, 20, 21]. We show below, that this constraint can be weakened if the lepton asymmetry, invoked in the sterile neutrino production, disappears at the specific stages. The reason is that the resonant production starts with particles of lower, than average, momenta and later proceeds to particles of higher momenta. Hence, if the lepton asymmetry disappears in the middle of the process, the final spectrum of the produced neutrinos turns colder than normally expected. Consequently, the average velocity of sterile neutrinos at the equality becomes lower, and then smaller sterile neutrino masses become consistent with the cosmic structure formation.
@@
We explain both options with an analysis of a simple Boltzmann equation describing the neutrino oscillations in plasma. We illustrate the expected results with the help of numerical codes dmpheno [22] and FortEPiaNO [23]. The first code evaluates the sterile neutrino production in a model, where all three neutrino asymmetries are taken into account. The second code traces the evolution of active neutrino asymmetries in the early Universe from the onset of active neutrino oscillations in the primordial plasma till the active neutrino decoupling.
@@
where $f_{\nu}(p,t)$ is neutrino $\nu$ distribution over 3-momentum $p$, $H(t)$ is the Hubble parameter, the expansion rate of the Universe, $\Gamma_{\nu_{\alpha}}$ is the neutrino $\nu_{\alpha}$ weak interaction rate in the plasma, which is of order $G_{F}^{2}T^{5}$ with Fermi constant $G_{F}$ and plasma temperature $T$. It is also assumed, that the weak interaction rate is much lower than the rate of oscillations between the active and sterile neutrinos, i.e.
@@
$V^{L}_{\alpha}=\sqrt{2}G_{F}\!\left[\!\left(1\!-\!2\sin^{2}\theta_{W}\right)\!\Delta n_{Q}-\frac{1}{2}\Delta n_{B}\!+\!\sum_{\beta}\left(\left(1+\delta_{\alpha\beta}\right)\Delta n_{\nu_{\beta}}\right)+\left(\!\delta_{\alpha\beta}\!-\!\frac{1}{2}\!+\!2\sin^{2}\theta_{W}\!\right)\!\Delta_{\beta}\right]$
with number densities of baryon charge $\Delta_{B}$, electric charge $\Delta n_{Q}$, neutral lepton asymmetries $\Delta n_{\nu_{\beta}}$ and charged lepton asymmetries $\Delta n_{\beta}$, see e.g. [24] for details. Hereafter we loosely use the same notation $p$ for neutrino momentum and neutrino energy, since all neutrinos are ultra-relativistic during the active period of sterile neutrino production, starting from temperature of several GeV to dozens of MeV.
Note that the production rate of sterile neutrinos grows proportionally to the weak interaction rate, see r.h.s. of eq. (1), since interaction in plasma collapses the neutrino wave function, which results in the appearance of sterile neutrino with probability (2). The weak rate is higher at higher temperature, but the transition probability (2) is highly suppressed both by damping (the second term in the denominator) and by thermal potential. Hence, generically the oscillation probability in plasma is lower than in vacuum, but for the lepton asymmetry potential which may cancel the mass term in the brackets and thus amplifies the transition similarly to the Mikheev–Smirnov–Wolfenstein effect in the Sun. The lepton asymmetry potential also decreases with temperature, $V^{L}_{\alpha}\propto T^{3}$, but slower than the thermal potential $V^{T}_{\alpha}\propto T^{5}$. Therefore, the oscillation probability is maximized when the quantity in brackets is close to zero,
@@
To analyze the expected spectrum it is useful to cast the Boltzmann equation in terms of variables $T$ and $y\equiv p/T$. The l.h.s. of the Boltzmann equation (1) transforms into $HT\partial f/\partial T$, while the resonance condition (3) may be written as follows
$\frac{A}{y}-B\cdot\Delta_{\alpha}\cdot T^{4}+C\cdot T^{6}y=0\,,$ (4)
where the positive parameters $A$, $B$, $C$ do not depend on the variables $y$ and $T$, and dimensionless parameter $\Delta_{\alpha}\equiv\Delta n_{\nu_{\alpha}}/s$ with entropy density $s$ characterizes the neutrino flavor asymmetry (the lepton flavor asymmetry is roughly proportional to it with the factor of about 3). Naturally, there are two solutions of the quadratic equation (4). Since the second and third terms exhibit strong dependence on temperature, one can consider the regimes of (relatively) large and small temperatures. In the first regime the solution is mostly saturated by the second and third terms in (4), so the resonance happens for the particles of momentum $p=yT$ with
$y\sim\Delta_{\alpha}\frac{B}{C}\,\frac{1}{T^{2}}\,.$ (5)
@@
$y\sim\frac{1}{\Delta_{\alpha}}\frac{A}{B}\,\frac{1}{T^{4}}\,.$ (6)
First, one observes, that in both cases (5), (6) the resonance production starts with the coldest particles. Since in (2) the damping term $\Gamma_{\nu_{\alpha}}^{2}\propto T^{10}$ suppresses transitions at high temperatures, it is the second solution (6) that matters in the production of sterile neutrinos. It begins to produce relatively cold sterile neutrinos and later proceeds to produce the warmer ones. This behavior is illustrated in Fig. 1
@@
Figure 1: Spectra of sterile neutrinos produced via mixing with electron neutrinos, $y=p/T$, at various plasma temperatures $T$ and various initial lepton asymmetries. The initial flavor neutrino asymmetries are equal to $(1,\,-0.5\,-0.5)\cdot N\cdot 10^{-3}$ for electron, muon and tau. The initial values of the scale factor $N$ used in calculations for the three plots are 1, 3,91391,\,3\,,91 , 3 , 9 from left to right. The dips in the spectrum are caused by the limited precision of the numerical calculations, which could be improved but at the cost of significantly increased computational time.
@@
by spectrum calculation performed with the help of routine dmpheno [22], the evolution starts from $T=4$ GeV and we adopt $m_{\nu_{s}}=7.11$ keV and $\sin^{2}2\theta=7\cdot 10^{-14}$. If the lepton asymmetry disappears, the second term in (4) vanishes, that ceases the sterile neutrino production. The final spectrum of the produced sterile neutrino is colder in this case, and hence lighter sterile neutrinos become consistent with the cosmic structure formation.
@@
Second, the resonance momentum becomes colder with higher asymmetry $\Delta_{\alpha}$. It is confirmed in Fig. 1 with numerical calculations for three different initial asymmetries. At the same temperature, the spectra are colder for higher asymmetries (from left to right on the plot) in full accordance with (6). Hence, exploiting higher asymmetry also helps to avoid the problems with cosmic structure formation and reopens a window of light sterile neutrinos. Moreover, with higher asymmetry the production starts later, reaching the epoch of active neutrino oscillations, $T\gtrsim 20$ MeV, which very efficiently cancel asymmetry between the neutrino flavors (see below) hence finishing the sterile neutrino production and consequently yielding generically colder spectrum.
@@
Alternatively, one may reduce the asymmetry and stop the production by involving some new ingredients. There is an issue here. Though, the presented in Fig. 1 spectra clearly show what can be achieved with instant disappearance of the lepton asymmetry, in realistic mechanisms the operation takes some time, and so the spectra are smoother than one could envisage from the plots. Actually, by reducing the asymmetry one shifts the most effective production to earlier stages but makes the production longer, and so the average momentum remains almost intact. We observed this behavior in a realistic situation with one more, but much heavier sterile neutrino component in the plasma. We assume, that these neutrinos are non-relativistic and their decay rate is about the Hubble parameter in the interesting epoch, $\Gamma_{d}=H(T_{d})$. The decay products are asymmetric in neutrino flavors and tend to reduce the neutrino asymmetries. This reduction changes the final spectra of the sterile neutrino very little, while we vary the temperature in the interval $26080<T_{d}<$ 260 MeV.
@@
One can see that the maxima of the final spectra shift towards the cold direction with earlier destruction of the lepton asymmetry (from possible decays of heavy fermions). There is no need to make the asymmetry equal to zero, just to reduce it significantly, e.g. for the right plot in Fig. 2 the final neutrino asymmetries are (1.93,0.04, $(-1.93,\,-0.04,\,3.75)\cdot 10^{-4}$. Reducing the asymmetry in this manner shifts the resonant production toward higher values of y, which results in a warmer spectrum. This effect competes with the cooling effect of the decays, which consequently limits the reduction in the mean momentum $p=yT$).
3. The larger lepton asymmetry yields the strongest amplification of the resonance production. One can check that at large asymmetry the amount of produced particles scales linearly with asymmetry. Thus, the larger the asymmetry, the lower the mixing angle $\theta$ is needed to provide the same amount of dark matter. However, when oscillations into sterile neutrino terminate, the net lepton asymmetry remains in plasma. In the literature, typically, it is supposed that the asymmetry resides mostly in one neutrino flavor. Then, the active neutrino oscillations in plasma, which begin at temperature of about 20 MeV, redistribute it over all three species. Then the electron neutrino asymmetry approaches one third of the initial asymmetry and must obey the upper limit inferred from observations of the helium abundance [16, 17]. It implies a lower limit on the mixing angle $\theta$ consistent with the hypothesis that sterile neutrinos, generated in this way, fully explain the dark matter component.
This limit may be significantly alleviated if the initial lepton asymmetry, used in the dark matter generation, is distributed between all three flavors in such a way that each flavor asymmetry is larger in value, but the total lepton asymmetry is close to zero. In such a situation, the active neutrino oscillations tend to reduce each of the flavor asymmetry, including the electron neutrino asymmetry, to zero, see Ref. [18].
@@
The evolution of neutrino asymmetry $L\equiv\sum_{\alpha}\Delta n_{\nu_{\alpha}}$ at the oscillation stage may be understood with simplified Boltzmann equation, assuming fast oscillations,
$\frac{dL}{dt}+3HL=-D\cdot\Gamma_{w}\,L\,,$
where dimensionless coefficient $D$ depends on the entries of neutrino mixing matrix. Replacing cosmic time $t$ with temperature $T$ and substituting the Hubble parameter as a function of temperature and Planck mass $M_{Pl}$, we find for the asymmetry-to-entropy ratio, $\Delta\equiv L/s$,
@@
where $F$ stands for some numerical factor. We observe that the exponential factor is determined by the temperature $T_{i}$ at the onset of oscillations, which occurs when the oscillation rate exceeds the weak interaction rate,
$\frac{\Delta m^{2}}{2T}\gtrsim G_{F}^{2}T^{5}\,,$
@@
$T_{i}\simeq\left(G_{F}^{2}\Delta m^{2}\right)^{1/6}\,.$
@@
$T_{\mu\tau}\approx 20\,\text{MeV}\,.$
Oscillations, which engage the electron neutrino asymmetry in this process, start at a lower temperature of about
$T_{e}\sim T_{\mu\tau}\times\left(\frac{\Delta m_{sol}^{2}}{\Delta m_{atm}^{2}}\right)^{1/6}\simeq 0.4\,T_{\mu\tau}\simeq 8\,\text{MeV}\,.$
Given the cubic dependence on temperature, it implies a factor of at least 100 smaller (as compared to the factor working for other two flavors) in the exponent (7), which eliminates the asymmetry in the electron neutrino component.
@@
Figure 3: Evolution of the neutrino asymmetries, electron neutrino (red lines), muon neutrino (green lines) and tau neutrino (blue lines) with temperature at the oscillation stage. The initial lepton asymmetries at 30 MeV are $(5.97,-3.01,-2.95)\cdot 10^{-4}$ (left panel) and $(87.9707,-46.9718,-35.6134)\cdot 10^{-4}$ (right panel). The neutrino oscillation parameters are taken from [16] for normal (solid lines) and inverse (dotted lines) ordering of the neutrino masses.
we illustrate this chain of reasoning by numerical calculations of sterile neutrino production performed with dmpheno code [22] for $m_{\nu_{s}}=7.11$ keV, mixing with electron neutrino, $\sin^{2}2\theta=7\cdot 10^{-14}$, and two different sets of initial neutrino flavor asymmetries, $(\Delta_{\nu_{e}},\Delta_{\nu_{\mu}},\Delta_{\nu_{\tau}})$, namely, $(2,-1,-1)\cdot 10^{-4}$ (left panel) and $(50,-25,-25)\cdot 10^{-4}$ (right panel) and subsequent evolution of the neutrino asymmetries at the stage of active neutrino oscillations, performed with FortEPiaNO code [23] and present estimates of the neutrino oscillation parameters [16]. The values of interest are the ratios of the neutrino chemical potentials to the temperature $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T$. They are related to the neutrino asymmetries $\Delta_{\nu_{\alpha}}$ used above as follows,

### v3

@@ Abstract
[…]
@@
However, $\theta$ may also be limited from below if the only source of sterile neutrino DM production is the sterile-active mixing  [^1] There are many models where other interactions are producing DM sterile neutrinos [11, 12, 13, 14, 15].. This is the case we study below.
@@
The lepton asymmetry may influence the later cosmological stages, in particular, primordial nucleosynthesis and recombination. Observations of primordial helium abundance favor standard cosmology and severely constrain [16, 17] the asymmetry in electron neutrinos at the moment of neutron freeze-out, which happens at the plasma temperature of about $T_{n}\approx 0.7$ MeV, some time after neutrino decoupling. As we show below, the lower bounds on mixing $\theta$, inferred from this constraint, can be significantly reduced in the case of a specific initial lepton asymmetry pattern, with large individual flavor asymmetries but a total lepton asymmetry close to zero. Then active neutrino oscillations, which start at temperature $T\lesssim 20$ MeV, very effectively cancel all flavor asymmetries in accordance with calculations of Refs. [18, 19]. We find that the cancellation is more efficient in the case of inverse ordering of neutrino masses, and CP-phase does not play a noticeable role. This allows the larger flavor lepton asymmetries at sterile neutrino production, hence lower mixing angles, consistent with observed helium abundance and anisotropy of the cosmic microwave background.
@@
The sterile neutrino mass $m_{\nu_{s}}$ is also bounded from below by observations of the cosmic structure formation. Being produced in oscillations of the active neutrino, the sterile neutrino inherits its momentum, which is generically of order of the plasma temperature. Hence, due to redshifting, a keV-scale sterile neutrinos exhibit velocities of order $10^{-3}$ at the onset of cosmic structure formation, at the epoch of equality between radiation and matter, when the plasma temperature is $T\approx 0.7$ eV. The dark matter candidates with this feature are called warm dark matter. The matter inhomogeneities at short distances are smoothed off by the free streaming of such dark matter particles. Observations of dwarf galaxies constrain the warm dark matter models with too high velocities, in particular, exclude too light sterile neutrinos, for recent studies see e.g. [20, 21, 22]. We show below, that this constraint can be weakened if the lepton asymmetry, invoked in the sterile neutrino production, disappears in the early Universe thus terminating the process. The reason is that the resonant production starts with particles of lower, than average, momenta and later proceeds to particles of higher momenta. Hence, if the lepton asymmetry disappears in the middle of the process, the final spectrum of the produced neutrinos turns colder than normally expected. Consequently, the average velocity of sterile neutrinos at the equality becomes lower, and then smaller sterile neutrino masses become consistent with the cosmic structure formation.
@@
We explain both options with an analysis of a simple Boltzmann equation describing the neutrino oscillations in plasma. We illustrate the expected results with the help of numerical codes dmpheno [23] and FortEPiaNO [24] ( which we slightly modified to suit our purposes). The first code evaluates the sterile neutrino production in a model, where all three neutrino asymmetries are taken into account. The second code traces the evolution of active neutrino asymmetries in the early Universe from the onset of active neutrino oscillations in the primordial plasma till the active neutrino decoupling.
@@
where $f_{\nu}(p,t)$ is neutrino $\nu$ distribution over 3-momentum $p$, $H(t)$ is the Hubble parameter, the expansion rate of the Universe, $\Gamma_{\nu_{\alpha}}$ is the neutrino $\nu_{\alpha}$ weak interaction rate in the plasma, which is of order $G_{F}^{2}T^{5}$ with Fermi constant $G_{F}$ and plasma temperature $T$. It is also assumed that the weak interaction rate is much lower than the rate of oscillations between the active and sterile neutrinos, i.e.
@@
$V^{L}_{\alpha}\!=\!\sqrt{2}G_{F}\!\left[\!\left(1\!-\!2\sin^{2}\theta_{W}\right)\!\Delta n_{Q}-\frac{1}{2}\Delta n_{B}\!+\!\sum_{\beta}\!\left(\left(1\!+\!\delta_{\alpha\beta}\right)\Delta n_{\nu_{\beta}}\right)+\left(\!\delta_{\alpha\beta}\!-\!\frac{1}{2}\!+\!2\sin^{2}\theta_{W}\!\right)\!\Delta n_{\beta}\right]$
with number densities of baryon charge $\Delta n_{B}$, electric charge $\Delta n_{Q}$, neutral lepton asymmetries $\Delta n_{\nu_{\beta}}$ and charged lepton asymmetries $\Delta n_{\beta}$, see e.g. [25] for details. Hereafter we loosely use the same notation $p$ for neutrino momentum and neutrino energy, since all neutrinos are ultra-relativistic during the active period of sterile neutrino production, starting from temperature of several GeV to dozens of MeV.
Note that the production rate of sterile neutrinos grows proportionally to the weak interaction rate, see r.h.s. of eq. (1), since interaction in plasma collapses the neutrino wave function, resulting in the appearance of a sterile neutrino with probability (2). The weak rate is higher at higher temperature, but the transition probability (2) is highly suppressed both by damping (the second term in the denominator) and by thermal potential. Hence, generically the oscillation probability in plasma is lower than in vacuum, but for the lepton asymmetry potential which may cancel the mass term in the brackets and thus amplifies the transition similarly to the Mikheev–Smirnov–Wolfenstein effect in the Sun. The lepton asymmetry potential also decreases with temperature, $V^{L}_{\alpha}\propto T^{3}$, but slower than the thermal potential, $V^{T}_{\alpha}\propto T^{5}$. Therefore, the oscillation probability is maximized when the quantity in brackets is close to zero,
@@
To analyze the expected spectrum analytically, it is useful to cast the Boltzmann equation in terms of variables $T$ and $y\equiv p/T$ and further neglect the decoupling of relativistic degrees of freedom from the plasma in the interesting epoch. Then l.h.s. of the Boltzmann equation (1) transforms into $HT\partial f/\partial T$, while the resonance condition (3) may be written as follows
$\frac{A}{y}\frac{m^{2}_{\nu_{s}}}{M_{W}^{2}}-B\frac{T^{4}}{M_{W}^{4}}\Delta_{\nu_{\alpha}}+C\frac{T^{6}}{M_{W}^{6}}y=0\,,$ (4)
where the positive dimensionless parameters $A$, $B$, $C$ (not necessary of order one) do not depend on the variables $y$, $T$ and can be directly extracted from the above formulas; dimensionless parameter $\Delta_{\nu_{\alpha}}\equiv\Delta n_{\nu_{\alpha}}/s$, with entropy density $s=2\pi^{2}h_{eff}T^{3}/45$[26] of plasma with $h_{eff}$ degrees of freedom, characterizes the neutrino flavor asymmetry (the lepton flavor asymmetry is roughly proportional to it with the factor of about 3). Naturally, there are two solutions of the quadratic equation (4). Since the second and third terms exhibit strong dependence on temperature, one can consider the regimes of (relatively) large and small temperatures. In the first regime the solution is mostly saturated by the second and third terms in (4), so the resonance happens for the particles of momentum $p=yT$ with
$y\sim y_{r1}\equiv\Delta_{\nu_{\alpha}}\frac{B}{C}\,\frac{M_{W}^{2}}{T^{2}}\,.$ (5)
@@
$y\sim y_{r2}\equiv\frac{1}{\Delta_{\nu_{\alpha}}}\frac{A}{B}\,\frac{m^{2}_{\nu_{s}}}{M_{W}^{2}}\frac{M_{W}^{4}}{T^{4}}\,.$ (6)
First, one observes that in both cases (5), (6) the resonance production starts with the coldest particles, because the resonance parameters $y_{r1}$, $y_{r2}$ grow with decreasing temperature. The ordering of the solutions (5), (6) depends on the dimensionless factor $\Delta_{\nu_{\alpha}}M_{W}/m_{\nu_{s}}$. Since in (2) the damping term $\Gamma_{\nu_{\alpha}}^{2}\propto G_{F}^{4}T^{10}$ suppresses transitions at high temperatures, it is the second solution (6) that matters in the production of sterile neutrinos. It begins to produce relatively cold sterile neutrinos and later proceeds to produce the warmer ones. This behavior is illustrated in Fig. 1
@@
Figure 1: Spectra of sterile neutrinos produced via mixing with electron neutrinos, $y=p/T$, at various plasma temperatures $T$ and various initial lepton asymmetries. The initial flavor neutrino asymmetries are equal to $(1,\,-0.5\,-0.5)\cdot N\cdot 10^{-3}$ for electron, muon and tau. The initial values of the scale factor $N$ used in calculations for the three plots are 1, 3,91,\,3\,,91 , 3 , 9 from left to right. The dips in the spectrum are caused by the limited precision of the numerical calculations, which could be improved but at the cost of significantly increased computational time.
@@
by the spectrum calculation performed with the help of routine dmpheno [23], the evolution starts from $T=4$ GeV and we consider the model with sterile neutrino mass $m_{\nu_{s}}=7.11$ keV and mixing with only electron neutrino described by $\sin^{2}2\theta=2.8\cdot 10^{-13}$. If the lepton asymmetry disappears, the second term in (4) vanishes, that ceases the sterile neutrino production. The final spectrum of the produced sterile neutrino is colder in this case, and hence lighter sterile neutrinos become consistent with the cosmic structure formation. Let us clarify that during the resonant oscillations some of the lepton asymmetry gets transferred to the sterile neutrinos. In the interesting region of tiny sterile-to-active mixing and relatively large initial lepton asymmetry, this transfer solely determines the final sterile neutrino abundance. However, numerically, the transfer is negligibly small, less than $10^{-4}$ of the initial lepton asymmetry. Moreover, we checked that we get the same numerical results while keeping the asymmetry constant at the evaluation. So, what we need above to cease the sterile neutrino production is a disappearance of the major part of the asymmetry.
@@
Second, the resonance momentum becomes colder with higher asymmetry $\Delta_{\nu_{\alpha}}$. It is confirmed in Fig. 1 with numerical calculations for three different initial asymmetries. At the same temperature, the spectra are colder for higher asymmetries (from left to right on the plot) in full accordance with (6). Hence, exploiting higher asymmetry also helps to avoid the problems with cosmic structure formation and reopens a window of light sterile neutrinos. Moreover, with higher asymmetry the production starts later, reaching the epoch of active neutrino oscillations, $T\gtrsim 20$ MeV, which very efficiently cancel asymmetry between the neutrino flavors (see below) hence finishing the sterile neutrino production and consequently yielding generically colder spectrum.
@@
Alternatively, one may reduce the asymmetry and stop the production by involving some new ingredients. There is an issue here. Though, the presented in Fig. 1 spectra clearly show what can be achieved with instant disappearance of the lepton asymmetry, in realistic mechanisms the operation takes some time, and so the spectra are smoother than one could envisage from the plots. Actually, by reducing the asymmetry one shifts the most effective production to earlier stages but makes the production longer, and so the average momentum remains almost intact. We observed this behavior in a realistic situation with one more, but much heavier sterile neutrino component in the plasma. We assume, that these neutrinos are non-relativistic and their decay rate is about the Hubble parameter in the interesting epoch, $\Gamma_{d}=H(T_{d})$. The decay products are asymmetric in neutrino flavors and tend to reduce the neutrino asymmetries. This reduction changes the final spectra of the sterile neutrino very little, while we vary the temperature in the interval $80<T_{d}<260$ MeV.
@@
[…]
3. A larger lepton asymmetry leads to stronger amplification of the resonance production. One can check that at large asymmetry the amount of produced particles scales linearly with asymmetry. Thus, the larger the asymmetry, the lower the mixing angle $\theta$ is needed to provide the same amount of dark matter. However, when oscillations into sterile neutrino terminate, the net lepton asymmetry remains in plasma. In the literature, typically, it is supposed that the asymmetry resides mostly in one neutrino flavor. Then, the active neutrino oscillations in plasma, which begin at temperature of about 20 MeV, redistribute it over all three species. Then the electron neutrino asymmetry approaches one third of the initial asymmetry and must obey the upper limit inferred from observations of the helium abundance [16, 17]. It implies a lower limit on the mixing angle $\theta$ consistent with the hypothesis that sterile neutrinos, generated in this way, fully explain the dark matter component.
This limit may be significantly alleviated if the initial lepton asymmetry, used in the dark matter generation, is distributed between all three flavors in such a way that each flavor asymmetry is larger in value, but the total lepton asymmetry is close to zero. In such a situation, the active neutrino oscillations tend to reduce each of the flavor asymmetry, including the electron neutrino asymmetry, to zero, see Refs. [18, 19].
@@
The redistribution of the asymmetry between the flavors at the fast oscillation stage is very natural, because the flavor changing processes are in equilibrium, while the total lepton number is conserved. It can be illustrated by considering oscillations between two species, say $\nu_{\tau}$ and $\nu_{\mu}$, where the redistribution process minimizes the difference $L_{\tau\mu}\equiv\Delta n_{\nu_{\tau}}-\Delta n_{\nu_{\mu}}$. The evolution of neutrino distributions over momenta $f_{\nu_{\alpha}}$ can be described with the simplified Boltzmann equations, e.g.
$\frac{\partial f_{\nu_{\tau}}}{\partial t}-Hp\frac{\partial f_{\nu_{\tau}}}{\partial p}=\frac{\Gamma_{w}(p)}{2}\langle P(\nu_{\mu}\to\nu_{\tau})\rangle\left(f_{\nu_{\mu}}-f_{\nu_{\tau}}\right)$ (7)
and similar for antineutrinos. Here only neutral currents mediate the weak interactions, and so the weak rate $\Gamma_{w}$ is the same for all equations. As in eq. (1) we assume that the oscillations are faster than the weak interactions, but also account for the inverse oscillation processes. We neglect all the processes not participating in the asymmetry redistribution, e.g. weak scatterings, which keep the distribution functions of the thermal form. The interesting asymmetry difference is determined by
$L_{\tau\mu}\propto\int d^{3}p\left(f_{\nu_{\tau}}-f_{\bar{\nu}_{\tau}}-f_{\nu_{\mu}}+f_{\bar{\nu}_{\mu}}\right)\,.$
Then forming a relevant combination by summing up the Boltzmann equations (7) and integrating over 3-momenta (one then replaces the momenta in $\Gamma_{w}$ by its thermal average) we obtain the equation for the asymmetry difference
$\frac{dL_{\tau\mu}}{dt}+3HL_{\tau\mu}=-D\cdot\Gamma_{w}(T)\,L_{\tau\mu}\,,$
where coefficient $D$ depends on the entries of neutrino mixing matrix.
This is an approximate equation but it captures all the main ingredients of the asymmetry distribution process in general case. Omitting the indices, replacing cosmic time $t$ with temperature $T$ and substituting the Hubble parameter as a function of temperature and Planck mass $M_{Pl}$, we find for the asymmetry-to-entropy ratio, $\Delta\equiv L/s$,
@@
where $F$ stands for some numerical factor. We observe that the exponential factor is determined by the temperature $T_{i}$ at the onset of oscillations, which occurs when the oscillation rate exceeds the weak interaction rate. More accurately, we are interested in the moment, when the vacuum term overcomes the matter term in the effective neutrino Hamiltonian. It implies, for average neutrino momenta,
$\frac{\Delta m^{2}}{2T}\gtrsim\varkappa G_{F}^{2}T^{5}\,,$
@@
$T_{i}\simeq\left(\varkappa^{-1}G_{F}^{-2}\Delta m^{2}\right)^{1/6}\,,$
with flavor-dependent numerical factor $\varkappa$.
@@
$T_{i,\mu\tau}\approx 20\,\text{MeV}\,.$
At this and lower temperatures the electron neutrinos interact in the plasma somewhat stronger than the muon and tau neutrinos, due to the presence of electron and positron pairs in the plasma, which yields $\varkappa_{e}/\varkappa_{\mu,\tau}\approx 4$[18]. Oscillations, which engage the electron neutrino asymmetry in this process, start at a lower temperature of about
$T_{i,e}\sim T_{i,\mu\tau}\times\left(\frac{\varkappa_{\tau}\Delta m_{sol}^{2}}{\varkappa_{e}\Delta m_{atm}^{2}}\right)^{1/6}\simeq 0.4\,T_{i,\mu\tau}\simeq 8\,\text{MeV}\,.$
The cubic dependence on temperature implies at least an order of magnitude smaller numerical factor (as compared to the factor working for other two flavors) in the exponent (8). Therefore, at the decoupling of neutrino from plasma, at the temperature of about $T_{d}\simeq 3$ MeV, the expected asymmetries would be
$\Delta_{\mu\tau}(T_{d})=\Delta_{\mu\tau}(T_{i,\mu\tau}){\mathrm{e}}^{-F_{\mu\tau}\cdot G_{F}^{2}M_{Pl}T_{i,\mu\tau}^{3}}\,,\;\;\;\Delta_{e}(T_{d})=\Delta_{e}(T_{i,e}){\mathrm{e}}^{-F_{e}\cdot G_{F}^{2}M_{Pl}T_{i,e}^{3}}\,.$
The oscillations involving electron neutrinos start later, and the limited time tends to make the asymmetry redistribution much less effective compared to the case of two other neutrinos. The final results also depend on the neutrino sector parameters (on the values of the parameters $F_{\mu\tau}$, $F_{e}$ in our illustrative example).
@@
Figure 3: Evolution of the neutrino asymmetries, electron neutrino (red lines), muon neutrino (green lines) and tau neutrino (blue lines) with plasma temperature, $x\equiv m_{e}/T$, at the oscillation stage. The initial lepton asymmetries at 30 MeV are $(5.97,-3.01,-2.95)\cdot 10^{-4}$ (left panel) and $(146.61,-78.29,-59.36)\cdot 10^{-4}$ (right panel). The neutrino oscillation parameters are taken from [16] for normal (solid lines) and inverse (dotted lines) ordering of the neutrino masses.
we illustrate this chain of reasoning by numerical calculations of sterile neutrino production performed with dmpheno code [23] for $m_{\nu_{s}}=7.11$ keV, mixing with electron neutrino, $\sin^{2}2\theta=2.8\cdot 10^{-13}$, and two different sets of initial neutrino flavor asymmetries, $(\Delta_{\nu_{e}},\Delta_{\nu_{\mu}},\Delta_{\nu_{\tau}})$, namely, $(2,-1,-1)\cdot 10^{-4}$ (left panel) and $(50,-25,-25)\cdot 10^{-4}$ (right panel) and subsequent evolution of the neutrino asymmetries at the stage of active neutrino oscillations, performed with FortEPiaNO code [24] and present estimates of the neutrino oscillation parameters [16]. The values of interest are the ratios of the neutrino chemical potentials to the temperature $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T$. They are related to the neutrino asymmetries $\Delta_{\nu_{\alpha}}$ used above as follows,